If you needed to make 100 mL of a 0.2 M fruit drink solution from the 1.0 M fruit drink solution, how would you do it? (Hint: Use MsVs = MdVd to find the amount of concentrated solution you need, then add water to reach 100 mL.)
1
Expert's answer
2019-05-28T06:21:18-0400

Using equation MsVs = MdVd (where Ms - molarity of stock solution, Vs - volume of stock solution, Md - molarity of diluted solution and Vd - volume of diluted solution) we can find Vs:

"V_s = \frac{M_dV_d}{M_s} = \frac {0.2\frac{mol}{L} \cdot 0.1 L}{1.0 \frac{mol}{L}}=0.02 L = 20 mL"

Thus, to prepare 100 mL of diluted solution we should measure 20 mL of concentrated stock solution and then add water to reach 100 mL.